# Sound velocity peak induced by the chiral partner in dense two-color QCD

### v1

@@ Abstract
Recently, the peak structure of the sound velocity was observed in the lattice simulation of two-color and two-flavor QCD at the finite quark chemical potential. The comparison with the chiral perturbation theory (ChPT) result was undertaken, however, the ChPT failed in reproducing the peak structure. In this study, to extend the ChPT framework, we incorporate contributions of the $\sigma$ meson, that is identified as the chiral partner of pions, on top of the low-energy pion dynamics by using the linear sigma model (LSM). Based on the LSM we derive analytic expressions of the thermodynamic quantities as well as the sound velocity within a mean-field approximation. As a result, we find that those quantities are provided by sums of the ChPT results and corrections, where the latter is characterized by a mass difference between the chiral partners, the $\sigma$ meson and pion. The chiral partner contributions are found to yield a peak in the sound velocity successfully. We furthermore show that the sound velocity peak emerges only when $m_{\sigma}>\sqrt{3}m_{\pi}$ and $\mu_{q}>m_{\pi}$, with $m_{\sigma(\pi)}$ and $\mu_{q}$ being the $\sigma$ meson (pion) mass and the quark chemical potential, respectively. The correlation between the sound velocity peak and the sign of the trace anomaly is also addressed.
@@ I Introduction
In QC2D, the diquarks, made of two quarks, form color-singlet baryons due to again the pseudoreality of $SU(2)_{c}$, and thus they obey the Bose-Einstein statistics. When the chemical potential $\mu_{q}$ reaches half the value of the pion mass, the ground-state diquark baryons become massless. Consequently, this gives rise to the Bose-Einstein condensation of diquarks, i.e., the diquark condensation, along with the creation of a baryonic matter, resulting in a transition to the baryon superfluid phase [6, 7]. In contrast to the large $\mu_{q}$, the hadronic phase is realized in the lower $\mu_{q}$, which is smoothly connected to the vacuum. Lattice simulations have been extensively performed to reveal various aspects in the cold dense QC2D such as phase structures, thermodynamics quantities and the hadron mass spectrum, across the phase transition [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
Recently, the $\mu_{q}$ dependence of the sound velocity has been evaluated in the QC2D lattice simulation [30]. The simulated result has indicated that in the baryon superfluid phase, the sound velocity monotonically grows from zero as $\mu_{q}$ increases, and soon it exceeds the conformal limit $c_{s}^{2}=1/3$. Meanwhile, at sufficiently large $\mu_{q}$ the relevant scale of the dense matter is solely provided by $\mu_{q}$, resulting in that the sound velocity eventually converges on $c_{s}^{2}=1/3$. Therefore, the observed excess implies the appearance of peak structures in the sound velocity.
@@
In sign-problem-free theories including QC2D, it has been pointed out that the confinement phase is favored in the cold dense regime, owing to the superfluidities [39, 40, 41]. These studies imply that the hadronic picture would be applicable even to the higher density region in the superfluid phase. Given this fact, an effective model approach based on hadronic degrees of freedom, which allows us to enter the superfluid phase, would be one of the reliable methods to elucidate properties of the sound velocity in dense QC2D.
As a conventional approach of such a hadron effective model, the chiral perturbation theory (ChPT) is frequently employed to make a comparison with lattice observations of cold dense QC2[6, 7, 42, 43, 41, 44]. In particular, the ChPT evaluation for the sound velocity within the mean-field approximation exhibits no explicit model parameter dependence [39, 13], and for this reason, the ChPT results serve as a robust benchmark of the low-energy QC2D in comparing with the lattice data. Indeed, the ChPT prediction of the sound velocity is in good agreement with the lattice data in the vicinity of the phase-transition point, however, it monotonically converges to $c_{s}^{2}=1$ at high density without showing any peak structures. Hence, there is a discrepancy between the QC2D lattice observation and the ChPT prediction slightly higher $\mu_{q}$. This is because the ChPT is constructed upon a low-energy expansion with respect to the Nambu-Goldstone (NG) bosons, and this model is adaptable for only the deep low-energy regime of QC2D. Thus, it is natural that the ChPT cannot apply to the high-dense regime, owing to the lack of higher excitations.
@@
This paper is organized as follows. In Sec. II we provide a brief review of the ChPT approach at a mean-field level, and present analytic expressions of the thermodynamics quantities including the sound velocity. Those quantities are also evaluated within the LSM approach in Sec. III, and the $\sigma$ meson contributions on top of the ChPT framework is clarified in this section. Then, in Sec. IV numerical demonstrations to visualize the $\sigma$ meson contributions are provided, and finally in Sec. V we conclude the present work.
@@ II ChPT in cold dense QC2D
In this study, we aim to delve into the understanding of the sound velocity from a perspective of the hadronic picture. In particular, we take into account hadron contributions absent in the ChPT framework and extend the hadronic result of the sound velocity beyond the low-energy regime governed by the ChPT. To facilitate our discussion, in this section, we provide a brief review of the ChPT which serves as a low-energy effective model for the lightest hadrons identified as the NG bosons associated with the chiral symmetry breaking. We also review the evaluation of the thermodynamic quantities, including the sound velocity, within the ChPT framework in finite quark chemical potentials at zero temperature.
@@ II.1 ChPT Lagrangian
In this subsection, we introduce the ChPT Lagrangian for two flavor QC2D which can access the baryonic matter.
In QC2D, two (anti)quarks can be bound together by the strong interaction to form color-singlet (anti)baryons in addition to mesons, carrying the quark number +2+2 (2-2). In contrast to the three-color QCD, those diquark baryons behave as bosonic particles sharing common properties with mesons due to the pseudoreal nature of color $SU(2)_{c}$ group. Hence, both the diquark baryons and mesons are treated on an equal footing in constructing the hadron effective model in QC2D. Besides, it is known that the pseudoreality allows us to extend $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry to the Pauli-Gursey $SU(4)$ symmetry, and in the low-energy region, this $SU(4)$ symmetry is spontaneously broken to the $Sp(4)$ symmetry [6, 7]. Therefore, in QC2D, low-energy effective models are constructed upon the symmetry-breaking pattern of $G=SU(4)\to H=Sp(4)$.
@@
Here, in two-flavor QC2D, the manifold of $G/H=SU(4)/Sp(4)$ is characterized by five degrees of freedom, and hence the superscript “ii” in Eq. (1) runs from $1$ to $5$. In terms of the hadrons, $\pi^{i=1,2,3}$ and $\pi^{i=4,5}$ denote the pseudoscalars and (anti)baryons seveing as the NG bosons, i.e., pions and positive-parity (anti)diquarks. The $4\times 4$ matrices $X^{i}$ are generators belonging to the Lie algebra of $G/H=SU(4)/Sp(4)$ given by
@@ II.2 ChPT in the baryon superfluid phase
As demonstrated by the lattice simulations in QC2D, the diquark baryonic matter is created, accompanied by the phase transition to the baryon superfluid phase, at which the chemical potential takes the half value of the vacuum pion mass, $\mu_{q}^{\rm cr}\equiv m_{\pi}/2$ [10, 21, 24]. These notable properties can also be described within the chiral effective models including the ChPT. In this subsection, based on the ChPT Lagrangian (7), we provide an analytical expression of the pressure in such a dense system. Moreover, we present key ingredients characterizing the baryonic matter: the quark-number density and the quark-number susceptibility.
@@
where $\beta$ lies in a range of $0\leq\beta\leq\pi/2$. By considering the matching with the underlying QC2D, we have the following correspondence between the quark bilinear condensates and $\beta$ in the VEV of $U$ [6, 7],
@@ II.3 Sound velocity in ChPT
Using the thermodynamic quantities derived in Sec. II.2, here, we present an analytical expression of the sound velocity in the ChPT within the mean-field approximation. In general, the sound velocity is defined by a derivative of the pressure $p$ with respect to the energy density $\epsilon$: $c_{s}^{2}=\partial p/\partial\epsilon$, along the isentropic curve. At zero temperature, the isentropic trajectory sticks to the $\mu_{q}$ axis, and then, the sound velocity can be simplified to the following concise expression [46]:
@@
The $\mu_{q}$ dependence of the sound velocity has been observed in the QC2D lattice simulation [30], indicating that the sound velocity departs from zero in the superfluid phase, and exceeds the conformal limit $c_{s}^{2}=1/3$. This conformal limit must be realized for sufficiently large $\mu_{q}$, since in such a dense matter the relevant scale is solely provided by $\mu_{q}$. Hence, the lattice result claims the appearance of the peak structure in the intermediate chemical potential regime in cold dense QC2D. This observation, however, contradicts the ChPT estimation (21). In what follows, we show that this lack of the peak structure in the ChPT framework can be resolved by including contributions from chiral partners of the NG bosons.
@@ III.1 LSM at mean-field approximation
where $X^{i=0}={\bm{1}}_{4\times 4}/(2\sqrt{2})$, and $X^{i}$ ($i=1$ - $5$) is defined in Eq. (2). The dynamical variables $\mathscr{S}^{i}$ and $\mathscr{P}^{i}$ are identified to the hadrons as follows: $\mathscr{S}^{i=0}$ is the iso-singlet scalar meson ($\sigma$ meson); $\mathscr{S}^{i=1,2,3}$ are the iso-triplet scalar mesons ($a_{0}$ mesons); $\mathscr{S}^{i=4,5}$ are the positive-parity (anti)baryons; $\mathscr{P}^{i=0}$ is the iso-singlet pseudoscalar meson ($\eta$ meson); $\mathscr{P}^{i=1,2,3}$ are the iso-triplet pseudoscalar meson (pions); $\mathscr{P}^{i=4,5}$ are the negative-parity (anti)baryons. It should be noted that $\Sigma$ has mass dimension +1+1 while $U$ is dimensionless, where these fields are related by $\Sigma\sim f_{\pi}U=f_{\pi}\xi^{2}E^{T}$.
Given that the difference between $\Sigma$ and $U$ is just the way of parametrization of the hadron field, $\Sigma$ also transforms under the Pauli-Gursey $SU(4)$ symmetry as
@@
where $m_{0}^{2}$ is a mass parameter, and $\lambda_{1}$ and $\lambda_{2}$ are dimensionless parameters controlling interactions among the hadrons. The last piece proportional to $m_{l}\bar{c}$ is responsible for the explicit breaking of the Pauli-Gursey $SU(4)$ symmetry which is inevitable to generate finite masses of the NG bosons. In Eq. (24), the $U(1)_{A}$ anomalous contributions are not introduced since the lattice results on the mass difference between $\eta$ and pion implies small effects from the $U(1)_{A}$ anomaly [45].
@@
The analytic expressions of $\sigma_{0}$ and $\Delta$ in Eq. (30) include explicit model parameters, which prevent us from achieving an intuitive picture of their $\mu_{q}$ dependences. For this reason, in what follows we try to translate $\sigma_{0}^{\rm(H)}$, $m_{0}^{2}$ and $\tilde{\lambda}$ into more physically transparent quantities. First, by evaluating the broken current associated with the broken $SU(4)/Sp(4)$ space, one can easily show that the mean field $\sigma_{0}^{\rm(H)}$ is connected to the pion decay constant as
$\sigma_{0}^{\rm(H)}=\sqrt{2}f_{\pi}.$ (33)
Next, from Eqs. (31) and (32), $m_{0}^{2}$ and $\tilde{\lambda}$ can be expressed in terms of the pion and $\sigma$ meson masses, which reads
@@
At the end of this subsection, we make comparisons between the ChPT and LSM results on $\mu_{q}$ dependences of the chiral and diquark condensates. From the matching result in Eq. (25) together with Eq. (35), one can derive the following relations with respect to the normalized chiral and diquark condensates in the superfluid phase:
@@
where the ChPT results on $\bar{\phi}$ and $\bar{d}$ are provided in Eq. (15). Thus, from Eq. (37), the normalized chiral condensate is found to share the common $\mu_{q}$ scaling in the superfluid phase. Meanwhile, the squared diquark condensate is divided into the ChPT result and an additional contribution $(\delta\bar{d}^{(\rm BS)})^{2}$. In particular, Eq. (38) indicates that the additional effects are quantified by the mass difference of the chiral partners, $\sigma$ meson and pion, and amplified as the chemical potential increases. Since $(\delta\bar{d}^{(\rm BS)})^{2}$ is proportional to $1/\delta\bar{m}_{\sigma-\pi}^{2}$, $\bar{d}^{(\rm BS)}_{\rm LSM}$ is reduced to $\bar{d}^{(\rm BS)}_{\rm ChPT}$ when the $\sigma$ meson mass is sufficiently large compared to the pion mass. This reduction is natural; such a heavy-mass limit is equivalent to integrating out the $\sigma$ meson from the LSM framework, which should converge on the ChPT. Hence, one can conclude that the $(\delta\bar{d}^{(\rm BS)})^{2}$ is induced as a correction beyond the lowest-energy dynamics governed by the NG bosons. Those clear properties of the LSM played by the $\sigma$ meson are also reflected to the thermodynamic quantities, as will be shown in the following analysis.

### v2

@@ Abstract
Recently, the peak structure of the sound velocity was observed in the lattice simulation of two-color and two-flavor QCD at the finite quark chemical potential. The comparison with the chiral perturbation theory (ChPT) result was undertaken; however, the ChPT failed in reproducing the peak structure. In this study, to extend the ChPT framework, we incorporate contributions of the $\sigma$ meson, that is identified as the chiral partner of pions, on top of the low-energy pion dynamics by using the linear sigma model (LSM). Based on the LSM we derive analytic expressions of the thermodynamic quantities as well as the sound velocity within a mean-field approximation. As a result, we find that those quantities are provided by sums of the ChPT results and corrections, where the latter is characterized by a mass difference between the chiral partners, the $\sigma$ meson and pion. The chiral partner contributions are found to yield a peak in the sound velocity successfully. We furthermore show that the sound velocity peak emerges only when $m_{\sigma}>\sqrt{3}m_{\pi}$ and $\mu_{q}>m_{\pi}$, with $m_{\sigma(\pi)}$ and $\mu_{q}$ being the $\sigma$ meson (pion) mass and the quark chemical potential, respectively. The correlation between the sound velocity peak and the sign of the trace anomaly is also addressed.
@@ I Introduction
In QC2D, the diquarks, made of two quarks, form color-singlet baryons due to again the pseudoreality of $SU(2)_{c}$, and thus they obey the Bose-Einstein statistics. When the chemical potential $\mu_{q}$ reaches half the value of the pion mass, the ground-state diquark baryons become massless. Consequently, this gives rise to the Bose-Einstein condensation of diquarks, i.e., the diquark condensation, along with the creation of a baryonic matter, resulting in a transition to the baryon superfluid phase [6, 7]. In contrast to the large $\mu_{q}$, the hadronic phase is realized in the lower $\mu_{q}$, which is smoothly connected to the vacuum. Lattice simulations have been extensively performed to reveal various aspects in the cold dense QC2D such as phase structures, thermodynamics quantities and the hadron mass spectrum, across the phase transition [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] (for a review, see [31]).
Recently, the $\mu_{q}$ dependence of the sound velocity has been evaluated in the QC2D lattice simulation [32]. The simulated result has indicated that in the baryon superfluid phase, the sound velocity monotonically grows from zero as $\mu_{q}$ increases, and soon it exceeds the conformal limit $c_{s}^{2}=1/3$. Meanwhile, at sufficiently large $\mu_{q}$ the relevant scale of the dense matter is solely provided by $\mu_{q}$, resulting in that the sound velocity eventually converges on $c_{s}^{2}=1/3$. Therefore, the observed excess implies the appearance of peak structures in the sound velocity. Note that the isospin chemical potential dependence of the sound velocity in the three-color QCD has also been investigated through lattice simulations [33, 34, 35, 36, 37], implying the presence of the sound velocity peak in the isospin asymmetric matter of the real-life QCD.
@@
In QC2D with large $\mu_{q}$, the gluon is unaffected by the Meissner effect [46, 47, 48] since the diquark condensate exists as a color singlet object. This leads to the gluon sector in QC2D being unscreened and implies that the highly dense QC2D system is expected to remain confined [49, 50, 51]. Indeed, the recent lattice simulation has shown that the Polyakov loop $L$ maintains small values even after the phase transition to the baryon superfluid phase: $L\ll 1$ [26, 24]. This indicates that QC2D would be in the confinement phase in a wide range of $\mu_{q}$ and the hadronic picture could be applicable across the baryon superfluid phase transition, although at the extremely high-density region the quark degrees of freedom become significant. The difficulty in clarifying the microscopic pictures of cold dense matter is related to the quark-hadron continuity [1]. In this study, we approach the baryonic matter from the low-energy perspective using hadronic models. These models allow us to continuously enter the baryon superfluid phase based on the hadronic picture, which would be one of the reliable methods to elucidate properties of the sound velocity from the low-energy viewpoint of dense QC2D.
As a conventional approach of such a hadron effective model, the chiral perturbation theory (ChPT) is frequently employed to make a comparison with lattice observations of cold dense QC2[6, 7, 52, 53, 51, 54]. In particular, the ChPT evaluation for the sound velocity within the mean-field approximation exhibits no explicit model parameter dependence [49, 13], and for this reason, the ChPT results serve as a robust benchmark of the low-energy QC2D in comparing with the lattice data. Indeed, the ChPT prediction of the sound velocity is in good agreement with the lattice data in the vicinity of the phase-transition point; however, it monotonically converges to $c_{s}^{2}=1$ at high density without showing any peak structures. Hence, there is a discrepancy between the QC2D lattice observation and the ChPT prediction of slightly higher $\mu_{q}$. This is because the ChPT is constructed upon a low-energy expansion with respect to the Nambu-Goldstone (NG) bosons, and this model is adaptable for only the deep low-energy regime of QC2D. Thus, it is natural that the ChPT cannot apply to the high-dense regime, owing to the lack of higher excitations.
@@
This paper is organized as follows. In Sec. II we provide a brief review of the ChPT approach at a mean-field level, and present analytic expressions of the thermodynamics quantities including the sound velocity. Those quantities are also evaluated within the LSM approach in Sec. III, and the $\sigma$ meson contributions on top of the ChPT framework is clarified in this section. Then, in Sec. IV, numerical demonstrations to visualize the $\sigma$ meson contributions are provided, and finally, in Sec. V, we conclude the present work.
@@ II ChPT in cold dense QC2D
In this study, we aim to delve into the understanding of the sound velocity from a perspective of the hadronic picture. In particular, we take into account hadron contributions absent in the ChPT framework and extend the hadronic result of the sound velocity beyond the low-energy regime governed by the ChPT. To facilitate our discussion, in this section, we provide a brief review of the ChPT which serves as a low-energy effective model for the lightest hadrons identified as the NG bosons associated with the chiral symmetry breaking. We also provide an overview of the well-established thermodynamic quantities in the ChPT. Following that, we review the evaluation of the sound velocity in finite quark chemical potentials at zero temperature.
@@ II.1 ChPT Lagrangian
In this subsection, we introduce the ChPT Lagrangian for two-flavor QC2D which can access the baryonic matter.
In QC2D, two (anti)quarks can be bound together by the strong interaction to form color-singlet (anti)baryons in addition to mesons, carrying the quark number $+2$ ($-2$). In contrast to the three-color QCD, those diquark baryons behave as bosonic particles sharing common properties with mesons due to the pseudoreal nature of color $SU(2)_{c}$ group. Hence, both the diquark baryons and mesons are treated on an equal footing in constructing the hadron effective model in QC2D. Besides, it is known that the pseudoreality allows us to extend $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry to the Pauli-Gürsey $SU(4)$ symmetry, and in the low-energy region, this $SU(4)$ symmetry is spontaneously broken to the $Sp(4)$ symmetry [6, 7]. Therefore, in QC2D, low-energy effective models are constructed upon the symmetry-breaking pattern of $G=SU(4)\to H=Sp(4)$.
@@
Here, in two-flavor QC2D, the manifold of $G/H=SU(4)/Sp(4)$ is characterized by five degrees of freedom, and hence the superscript “$i$” in Eq. (1) runs from $1$ to $5$. In terms of the hadrons, $\pi^{i=1,2,3}$ and $\pi^{i=4,5}$ denote the pseudoscalars and (anti)baryons serving as the NG bosons, i.e., pions and positive-parity (anti)diquarks. The $4\times 4$ matrices $X^{i}$ are generators belonging to the Lie algebra of $G/H=SU(4)/Sp(4)$ given by
@@ II.2 ChPT in the baryon superfluid phase
As demonstrated by the lattice simulations in QC2D, the diquark baryonic matter is created, accompanied by the phase transition to the baryon superfluid phase, at which the chemical potential takes the half value of the vacuum pion mass, $\mu_{q}^{\rm cr}\equiv m_{\pi}/2$ [10, 21, 24]. These notable properties can also be described within the chiral effective models including the ChPT. In this subsection, we recapitulate the well-known analytic expressions of the thermodynamic quantities in the ChPT [49, 13]: the pressure, the quark-number density, the quark-number susceptibility and the energy density.
@@
where $\beta$ lies in a range of $20\leq\beta\leq\pi/$ 2. By considering the matching with the underlying QC2D, we have the following correspondence between the quark bilinear condensates and $\beta$ in the VEV of $U$ [6, 7],
@@ II.3 Sound velocity in ChPT
Using the thermodynamic quantities shown in Sec. II.2, one can obtain an analytic expression of the sound velocity in the ChPT within the mean-field approximation. In general, the sound velocity is defined by a derivative of the pressure $p$ with respect to the energy density $\epsilon$: $c_{s}^{2}=\partial p/\partial\epsilon$, along the isentropic curve. At zero temperature, the isentropic trajectory sticks to the $\mu_{q}$ axis, and then, the sound velocity can be simplified to the following concise expression [56]:
@@
The $\mu_{q}$ dependence of the sound velocity has been observed in the QC2D lattice simulation [32], indicating that the sound velocity departs from zero in the superfluid phase and exceeds the conformal limit $c_{s}^{2}=1/3$. This conformal limit must be realized for sufficiently large $\mu_{q}$, since in such a dense matter the relevant scale is solely provided by $\mu_{q}$. Hence, the lattice result claims the appearance of the peak structure in the intermediate chemical potential regime in cold dense QC2D. This observation, however, contradicts the ChPT estimation (21). In what follows, we show that this lack of the peak structure in the ChPT framework can be resolved by including contributions from chiral partners of the NG bosons.
@@ III.1 LSM at mean-field approximation
where $X^{i=0}={\bm{1}}_{4\times 4}/(2\sqrt{2})$, and $X^{i}$ ($i=1$ $5$) is defined in Eq. (2). The dynamical variables $\mathscr{S}^{i}$ and $\mathscr{P}^{i}$ are identified to the hadrons as follows: $\mathscr{S}^{i=0}$ is the isosinglet scalar meson ($\sigma$ meson); $\mathscr{S}^{i=1,2,3}$ are the isotriplet scalar mesons ($a_{0}$ mesons); $\mathscr{S}^{i=4,5}$ are the positive-parity (anti)baryons; $\mathscr{P}^{i=0}$ is the isosinglet pseudoscalar meson ($\eta$ meson); $\mathscr{P}^{i=1,2,3}$ are the isotriplet pseudoscalar meson (pions); $\mathscr{P}^{i=4,5}$ are the negative-parity (anti)baryons. It should be noted that $\Sigma$ has mass dimension $+1$ while $U$ is dimensionless, where these fields are related by $\Sigma\sim f_{\pi}U=f_{\pi}\xi^{2}E^{T}$.
Given that the difference between $\Sigma$ and $U$ is just the way of parametrization of the hadron field, $\Sigma$ also transforms under the Pauli-Gürsey $SU(4)$ symmetry as
@@
where $m_{0}^{2}$ is a mass parameter, and $\lambda_{1}$ and $\lambda_{2}$ are dimensionless parameters controlling interactions among the hadrons. The last piece proportional to $m_{l}\bar{c}$ is responsible for the explicit breaking of the Pauli-Gürsey $SU(4)$ symmetry which is inevitable to generate finite masses of the NG bosons. In Eq. (24), the $U(1)_{A}$ anomalous contributions are not introduced since the lattice results on the mass difference between $\eta$ and pion imply small effects from the $U(1)_{A}$ anomaly [55].
@@
The analytic expressions of $\sigma_{0}$ and $\Delta$ in Eq. (30) include explicit model parameters, which prevent us from achieving an intuitive picture of their $\mu_{q}$ dependencies. For this reason, in what follows we try to translate $\sigma_{0}^{\rm(H)}$, $m_{0}^{2}$ and $\tilde{\lambda}$ into more physically transparent quantities. First, by evaluating the broken current associated with the broken $SU(4)/Sp(4)$ space, one can easily show that the mean field $\sigma_{0}^{\rm(H)}$ is connected to the pion decay constant as
$\sigma_{0}^{\rm(H)}=\sqrt{2}f_{\pi},$ (33)
which is derived in the Appendix. Next, from Eqs. (31) and (32), $m_{0}^{2}$ and $\tilde{\lambda}$ can be expressed in terms of the pion and $\sigma$ meson masses, which reads
@@
At the end of this subsection, we make comparisons between the ChPT and LSM results on $\mu_{q}$ dependencies of the chiral and diquark condensates. From the matching result in Eq. (25) together with Eq. (35), one can derive the following relations with respect to the normalized chiral and diquark condensates in the superfluid phase:
@@
where the ChPT results on $\bar{\phi}$ and $\bar{d}$ are provided in Eq. (15). Thus, from Eq. (37), the normalized chiral condensate is found to share the common $\mu_{q}$ scaling in the superfluid phase. Meanwhile, the squared diquark condensate is divided into the ChPT result and an additional contribution $(\delta\bar{d}^{(\rm BS)})^{2}$. In particular, Eq. (38) indicates that the additional effects are quantified by the mass difference of the chiral partners, $\sigma$ meson and pion, and amplified as the chemical potential increases. Since $(\delta\bar{d}^{(\rm BS)})^{2}$ is proportional to $1/\delta\bar{m}_{\sigma-\pi}^{2}$, $\bar{d}^{(\rm BS)}_{\rm LSM}$ is reduced to $\bar{d}^{(\rm BS)}_{\rm ChPT}$ when the $\sigma$ meson mass is sufficiently large compared to the pion mass. This reduction is natural; such a heavy-mass limit is equivalent to integrating out the $\sigma$ meson from the LSM framework, which should converge on the ChPT. Hence, one can conclude that the $(\delta\bar{d}^{(\rm BS)})^{2}$ is induced as a correction beyond the lowest-energy dynamics governed by the NG bosons. Those clear properties of the LSM played by the $\sigma$ meson are also reflected to the thermodynamic quantities, as is shown in the following analysis.